# ARMFlow: AutoRegressive MeanFlow for Online 3D Human Reaction Generation

### v1

@@ Abstract
3D human reaction generation faces three main challenges: (1) high motion fidelity, (2) real-time inference, and (3) autoregressive adaptability for online scenarios. Existing methods fail to meet all three simultaneously. We propose ARMFlow, a MeanFlow-based autoregressive framework that models temporal dependencies between actor and reactor motions. It consists of a causal context encoder and an MLP-based velocity predictor. We introduce Bootstrap Contextual Encoding (BSCE) in training, encoding generated history instead of the ground-truth ones, to alleviate error accumulation in autoregressive generation. We further introduce the offline variant ReMFlow, achieving state-of-the-art performance with the fastest inference among offline methods. Our ARMFlow addresses key limitations of online settings by: (1) enhancing semantic alignment via a global contextual encoder; (2) achieving high accuracy and low latency in a single-step inference; and (3) reducing accumulated errors through BSCE. Our single-step online generation surpasses existing online methods on InterHuman and InterX by over 40% in FID, while matching offline state-of-the-art performance despite using only partial sequence conditions. Code is available in the supplementary.
@@ 1 Introduction
Recent advances in generative modeling have led to remarkable progress in 3D human motion generation, covering a wide spectrum of tasks and methodologies. These include text-guided motion synthesis [a2m, t2mgpt, temos, mdm, mld, guo2024momask, motiondiffuse, attt2m, fgt2m], human–object interaction [interdiff, hoidiff, chois, ardhoi, cghoi], scene-conditioned motion generation, and multi-person interaction modeling. Among these, one particularly distinctive task is human reaction generation, which focuses on producing reactive human behaviors in response to other agents or stimuli. This task holds immediate practical potential in human–robot interaction, augmented, and virtual realities.
Figure 1: Our method only processes a single inference in each real-time step for online reaction generation, compared to the SOTA methods ReGenNet (35-78 ms), and CAMDM (45 ms). The text description is from the InterX [interx] dataset.
@@
Existing approaches have made attempts to address these challenges, yet fundamental limitations remain. To enable online generation, autoregressive architectures are essential, as they inherently model temporal dependencies between past and future motions. Methods such as CAMDM [camdm] and HumanX [ji2025humanx] adopt autoregressive diffusion models, where a fixed-length historical window serves as the conditioning context for current denoising steps. However, this design faces two key limitations: (1) the fixed context window hinders scalability and leads to inevitable information loss, causing semantic drift over long sequences; and (2) although accelerated samplers such as DDIM have been adopted, they still require multiple denoising steps (typically $\geq$8), which is computationally demanding, especially under fine-grained temporal resolutions.
To address the aforementioned three challenges and two key limitations of prior works, we propose ARMFlow, a scalable autoregressive architecture capable of generating high-fidelity human reactions in a single-step inference. Our method is built upon the recently proposed MeanFlow [geng2025mean] paradigm, which enables one-step generation as opposed to multi-step denoising or iterative integration required by diffusion or traditional flow-based models. This is the first work that leverages MeanFlow for human motion generation, demonstrating the fastest inference speed while maintaining superior motion realism compared to existing approaches. To overcome the limitations of fixed-length contextual windows commonly used in prior autoregressive models, we introduce a causal context encoder that encodes the entire motion history with causal masking. This design prevents the loss of information beyond the context window and preserves global temporal semantics, ensuring coherent long-term motion generation.
@@ 2 Related Works
3D human motion generation is currently an active area of research, encompassing tasks such as Text-to-Motion [a2m, temos, mdm, mld, ACTOR, t2mgpt, guo2024momask, attt2m, fgt2m], Human-Object Interaction [hoidiff, cghoi, chois, ardhoi, interdiff, physhoi], to Human-Human Interaction [intergen, in2in, wang2025timotion, intermask] generation. However, Action-Reaction Generation is still an understudied area. Tab. 1 lists current SOTA models for reaction generation.
@@
InterMask [intermask]
InterGen [intergen]
MARRS [wang2025marrs]
CAMDM [camdm]
ReGenNet [xu2024regennet]
HumanX [ji2025humanx]
@@
interformer first proposed Interformer, a Transformer-based [vaswani2017attention] reaction synthesis model, but due to the use of traditional autoencoders to predict current actions, its prediction accuracy is less than ideal. xu2024regennet introduced the first online Transformer-decoder-based diffusion model, aiming to achieve both efficiency and accuracy in diffusion via DDIM [ddim]. Although this method performs online generation, it inherently lacks autoregression (AR) (i.e., generating the current action based on previously generated content), leading to temporal inconsistency. ji2025humanx addressed the autoregressive online generation issue by conditioning on a fixed window of previous context to guide the generation of the next timestep’s action. However, a fixed-window design imposes significant scalability constraints: it cannot be adapted to finer temporal resolutions, and approaches such as [camdm] are unable to perform inference from scratch without any prior context. Moreover, when processing long sequences, the limited receptive field fails to capture distant dependencies, leading to cumulative error and lack of global contextual understanding.
@@
have been extensively explored, particularly for image generation. Some approaches tokenize continuous representations into discrete token spaces in an LLM-style manner [tian2024var, lfqvae, t2mgpt, maskgit], while others adopt diffusion-based methods [ardiff, ARDM, armd, aamdm] in an autoregressive fashion. Another line of research employs autoregressive flow-based architectures [tokenflow, nextstep]. Although autoregressive models are capable of capturing strong dependencies within data distributions, a core challenge lies in mitigating error accumulation that naturally occurs during sequential inference. T2M-GPT [t2mgpt] addresses this issue by introducing random token corruption during training, forcing the model to correct corrupted past contexts. Subsequently, tian2024var proposed a hierarchical data representation to alleviate cumulative errors. However, these techniques are not directly applicable to conventional autoregressive models—especially diffusion-based architectures with fixed contextual windows [ARDM, armd, aamdm]. To overcome this limitation, ardiff proposed a general framework combining an autoregressive context encoder with a lightweight MLP denoiser, leveraging MAE-style random masking [MAE] to reduce error propagation along the autoregressive direction. Nevertheless, such random masking schemes are less suitable for specially ordered or real-time generation tasks, where temporal consistency is essential. As a result, enhancing the robustness of real-time autoregressive models remains unsolved.
@@ 3 Preliminary
MeanFlow [geng2025mean] is a recently proposed generative paradigm that aims to produce high-fidelity samples with only a single inference step. Unlike traditional Flow Matching (FM) [flowmatching, rectifiedflow] approaches, which learn the instantaneous velocity field by estimating the continuous trajectory between data and noise distributions through numerous integration steps, MeanFlow instead models the mean velocity over the entire trajectory. Specifically, instead of computing the instantaneous flow $\mathbf{v}(x,t)$ at each time step, MeanFlow evaluates the average transport velocity between any two points $r$ and $t$ in the field. This simplification allows the model to generate samples by a single-step integration of the mean field from $r=0$ (noise) to $t=1$ (data), substantially reducing inference time while maintaining strong generative quality. Our method builds upon MeanFlow, which predicts an average velocity field that bridges a Gaussian noise distribution and the target motion distribution, enabling efficient offline reaction generation with controllable conditioning and single-step inference. Given an instantaneous velocity field $v(z_{\tau},\tau)$, the average velocity between timesteps $r$ and $t$ is defined as:
@@
where $\mathrm{sg}(\cdot)$ denotes the stop-gradient operator to stabilize optimization. The Jacobian–vector products (JVPs) for $\partial_{z}u_{\theta}$ and $\partial_{t}u_{\theta}$ are computed via automatic differentiation, avoiding explicit Jacobian construction and reducing memory overhead while maintaining accurate gradient signals. Unlike conventional flow-based methods that uniformly sample timestep pairs, we adopt a biased sampling strategy where a proportion of pairs satisfy $r=t$. This allocation strengthens the learning of the instantaneous velocity $v(z_{t},t)$, improves convergence, and enhances fidelity for high-frequency motion components, which are critical for reactive behaviors. A key advantage of MeanFlow is single-step inference, eliminating iterative refinement:
$\hat{z}_{0}=\epsilon-u_{\theta}(\epsilon,0,1),\qquad\epsilon\sim\mathcal{N}(0,I),$ (4)
which drastically reduces inference cost compared to multi-step diffusion or flow-based models, making it suitable for large-scale generation.
To improve controllability without additional inference overhead, geng2025mean incorporate classifier-free guidance (CFG) during training rather than post-hoc blending by constructing a Ground-Truth Field. The modified objective becomes:
@@ 4 Method
Given a text description texttext and an actor motion sequence $x_{a}\in\mathbb{R}^{T,D}$, where $T$ denotes the number of frames and $D$ the pose dimension, our goal is to synthesize a realistic reactor motion $x_{b}\in\mathbb{R}^{T,D}$. The generation can occur in two modes: sequential (online), where frames are predicted progressively, or offline, where the entire sequence is generated in a single pass. To achieve this, we first compress actor and reactor motions into a shared latent space using the CNN-VAE described in Section 4.1, which benefits both offline and online reaction generation. Building on this representation, we introduce ReMFlow (Section 4.2) for offline generation and ARMFlow (Section 4.3) for online generation, both leveraging the MeanFlow framework and a DiT-based backbone [dit]. Finally, we propose the Bootstrap Context Encoding (BCSE), which is inherently suited for ARMFlow and mitigates error accumulation during autoregression.
@@ 4.1 CNN-VAE for Motion Compression
To enhance generation efficiency and model robustness, we adopt a 1D-CNN VAE following the T2M-GPT [t2mgpt] architecture to compress actor–reactor motion sequences $\mathbf{x}=\{\mathbf{x}_{a},\mathbf{x}_{b}\}$ into two temporal latent representations $\{\mathbf{a},\mathbf{b}\}$. This compression serves two purposes: (1) It reduces the dimensionality of motion data for efficient transmission and computation, and (2) provides a structured latent space that facilitates autoregressive modeling.
Unlike canonicalized actor motion, reactor motion is highly dynamic and lacks a standardized representation, making discrete quantization prone to severe accuracy degradation. To address this, we employ continuous latent tokens with KL-divergence regularization instead of discrete codes. This design embeds motion into a compact yet expressive latent space, accelerating convergence and improving reconstruction fidelity. Furthermore, the causal nature of 1D-CNN ensures that temporal dependencies are preserved, enabling efficient disentanglement of latent tokens in online generation. This property significantly enhances flexibility when handling sequential contexts.
The VAE objective is defined as:
@@
where $q(z|x)$ is the encoder distribution, $p(x|z)$ is the decoder likelihood, and $p(z)$ is the Gaussian prior. To facilitate high-fidelity VAE reconstruction, we employ inverse kinematic (IK) loss for explicit joint position and velocity loss for motion smoothness.
@@ 4.2 ReMFlow for Offline Reaction Generation
We implement ReMFlow for offline generation using a DiT-based encoder with multimodal conditioning. Text prompts are encoded via a CLIP text encoder [CLIP] and fused with timestep embeddings to form a global condition vector, which is injected into DiT through adaptive layer normalization (AdaLN) to modulate intermediate activations. Actor tokens $a_{i}$ are concatenated with interpolated reactor tokens $b_{i}^{t}$, encoded from a CNN-based VAE to capture local appearance and dynamics of the reactor. These tokens are aligned along the last dimension, summed with positional encodings, and processed by DiT to predict average velocities $\mathbf{u}=\{u_{i}\}_{i=1}^{m}$ over tokenized spatiotemporal patches. To enable unconditional learning and support CFG, a proportion of text and actor tokens are replaced with null tokens $\varnothing$ during training, which regularizes the model and stabilizes optimization under varying conditioning strengths.
@@ 4.3 ARMFlow for Online Generation
ARMFlow overcomes the limitations of fixed-window architectures by enabling encoding from scratch and retaining all past information, thereby preserving global semantics. As illustrated in Fig. 2, the architecture follows the spirit of MAR [ardiff], comprising a DiT-based context encoder and a lightweight MLP velocity predictor. The former encodes historical motion and text semantics and supplies them as partial conditions to the MLP predictor.
@@
At inference time, the start token $\langle\text{sos}\rangle$ is first encoded as the initial history. A Gaussian noise token $\epsilon_{1}$ is then paired with the current actor token $a_{1}$ and fed into the MLP with (r=0,t=1)(r{=}0,\,t{=}1) to predict the average velocity $\hat{u}_{1}$. The current reactor token is obtained by $b_{1}=\epsilon_{1}-\hat{u}_{1}$ according to Eq.4. The predicted reactor token is cached together with the corresponding actor token to update the history, and the process is iterated autoregressively until actor tokens end.
@@ 4.4 Bootstrap Contextual Encoding
Drift over long sequences is an inherent limitation of autoregressive models. HumanX [ji2025humanx] addresses this through a history-rollout training strategy, where the ground-truth (GT) history condition is gradually replaced by the model’s generated history. However, this approach has three main drawbacks. First, to stabilize training, HumanX gradually subtracts the GT history from the generated ones, slowing the convergence. Second, as the model converges, the generated history becomes increasingly similar to the GT, leading to insufficient self-augmentation. Third, HumanX only replaces the reactor history while keeping the actor unchanged, which lead to overfitting to the actor’s motion.
@@
InterFormer [interformer] $4.871^{\pm.049}$ ${0.302}^{\pm.004}$ ${0.457}^{\pm.004}$ ${0.542}^{\pm.005}$ ${3.845}^{\pm.001}$ $7.482^{\pm.045}$ ${0.254}^{\pm.029}$
CAMDM [camdm] $\underline{4.000}^{\pm.046}$ ${0.335}^{\pm.005}$ ${0.492}^{\pm.005}$ ${0.587}^{\pm.005}$ ${3.828}^{\pm.001}$ $\underline{7.547}^{\pm.025}$ $\textbf{1.581}^{\pm.026}$
ReGenNet [xu2024regennet] ${4.176}^{\pm.085}$ $\underline{0.355}^{\pm.005}$ $\underline{0.508}^{\pm.005}$ $\underline{0.600}^{\pm.004}$ $\underline{3.817}^{\pm.001}$ ${7.480}^{\pm.033}$ $\underline{0.442}^{\pm.012}$
@@
InterFormer [interformer] ${0.304}^{\pm.009}$ ${0.301}^{\pm.003}$ ${0.469}^{\pm.003}$ ${0.571}^{\pm.002}$ ${4.604}^{\pm.009}$ ${8.579}^{\pm.061}$ ${0.289}^{\pm.009}$
CAMDM [camdm] ${0.429}^{\pm.011}$ ${0.312}^{\pm.004}$ ${0.480}^{\pm.003}$ ${0.587}^{\pm.003}$ ${4.468}^{\pm.020}$ ${8.467}^{\pm.072}$ $\textbf{1.460}^{\pm.027}$
ReGenNet [xu2024regennet] $\underline{0.071}^{\pm.003}$ $\underline{0.402}^{\pm.005}$ $\underline{0.584}^{\pm.004}$ $\underline{0.690}^{\pm.004}$ $\underline{3.843}^{\pm.011}$ $\underline{9.011}^{\pm.053}$ ${0.738}^{\pm.021}$

### v2

@@ Abstract
3D human reaction generation faces three main challenges: (1) high motion fidelity, (2) real-time inference, and (3) autoregressive adaptability for online scenarios. Existing methods fail to meet all three simultaneously. We propose ARMFlow, a MeanFlow-based autoregressive framework that models temporal dependencies between actor and reactor motions. It consists of a causal context encoder and an MLP-based velocity predictor. We introduce Bootstrap Contextual Encoding (BSCE) in training, encoding generated history instead of the ground-truth ones, to alleviate error accumulation in autoregressive generation. We further introduce the offline variant ReMFlow, achieving state-of-the-art performance with the fastest inference among offline methods. Our ARMFlow addresses key limitations of online settings by: (1) enhancing semantic alignment via a global contextual encoder; (2) achieving high accuracy and low latency in a single-step inference; and (3) reducing accumulated errors through BSCE. Our single-step online generation surpasses existing online methods on InterHuman and InterX by about 30% in FID, while matching offline state-of-the-art performance despite using only partial sequence conditions. The official implementation is publicly available at: https://github.com/ZenGengChin/armflow_official
@@ 1 Introduction
Recent advances in generative modeling have led to remarkable progress in 3D human motion generation, covering a wide spectrum of tasks and methodologies. These include text-guided motion synthesis [12, 47, 32, 36, 4, 11, 48, 49, 41], human–object interaction [45, 30, 18, 9, 6], scene-conditioned motion generation, and multi-person interaction modeling. Among these, one particularly distinctive task is human reaction generation, which focuses on producing reactive human behaviors in response to other agents or stimuli. This task holds immediate practical potential in human–robot interaction, augmented, and virtual realities.
Figure 1: Our method only processes a single inference in each real-time step for online reaction generation, compared to the SOTA methods ReGenNet (35-78 ms), and CAMDM (45 ms). The text description is from the InterX [43] dataset.
@@
Existing approaches have made attempts to address these challenges, yet fundamental limitations remain. To enable online generation, autoregressive architectures are essential, as they inherently model temporal dependencies between past and future motions. Methods such as CAMDM [3] and HumanX [17] adopt autoregressive diffusion models, where a fixed-length historical window serves as the conditioning context for current denoising steps. However, this design faces two key limitations: (1) the fixed context window hinders scalability and leads to inevitable information loss, causing semantic drift over long sequences; and (2) although accelerated samplers such as DDIM have been adopted, they still require multiple denoising steps (typically $\geq$8), which is computationally demanding, especially under fine-grained temporal resolutions.
To address the aforementioned three challenges and two key limitations of prior works, we propose ARMFlow, a scalable autoregressive architecture capable of generating high-fidelity human reactions in a single-step inference. Our method is built upon the recently proposed MeanFlow [8] paradigm, which enables one-step generation as opposed to multi-step denoising or iterative integration required by diffusion [23, 9] or traditional flow-based models. This is the first work that leverages MeanFlow for human motion generation, demonstrating the fastest inference speed while maintaining superior motion realism compared to existing approaches. To overcome the limitations of fixed-length contextual windows commonly used in prior autoregressive models, we introduce a causal context encoder that encodes the entire motion history with causal masking. This design prevents the loss of information beyond the context window and preserves global temporal semantics, ensuring coherent long-term motion generation.
@@ 2 Related Works
3D human motion generation is currently an active area of research, encompassing tasks such as Text-to-Motion [12, 32, 36, 4, 31, 47, 11, 49, 41], Human-Object Interaction [30, 6, 18, 9, 45, 42], to Human-Human Interaction [21, 33, 39, 16] generation. However, Action-Reaction Generation is still an understudied area. Tab. 1 lists current SOTA models for reaction generation.
@@
InterMask [16]
InterGen [21]
MARRS [40]
CAMDM [3]
ReGenNet [44]
HumanX [17]
@@
Chopin et al. [5] first proposed Interformer, a Transformer-based [38] reaction synthesis model, but due to the use of traditional autoencoders to predict current actions, its prediction accuracy is less than ideal. Xu et al. [44] introduced the first online Transformer-decoder-based diffusion model, aiming to achieve both efficiency and accuracy in diffusion via DDIM [34]. Although this method performs online generation, it inherently lacks autoregression (AR) (i.e., generating the current action based on previously generated content), leading to temporal inconsistency. Ji et al. [17] addressed the autoregressive online generation issue by conditioning on a fixed window of previous context to guide the generation of the next timestep’s action. However, a fixed-window design imposes significant scalability constraints: it cannot be adapted to finer temporal resolutions, and approaches such as [3] are unable to perform inference from scratch without any prior context. Moreover, when processing long sequences, the limited receptive field fails to capture distant dependencies, leading to cumulative error and a lack of global contextual understanding. R2R [1] aims to solve this by introducing full-history encoding, but its 50-step inference limits its efficiency.
@@
have been extensively explored, particularly for image generation. Some approaches tokenize continuous representations into discrete token spaces in an LLM-style manner [37, 46, 47, 2], while others adopt diffusion-based methods [19, 15, 13, 20] in an autoregressive fashion. Another line of research employs autoregressive flow-based architectures [10, 35]. Although autoregressive models are capable of capturing strong dependencies within data distributions, a core challenge lies in mitigating error accumulation that naturally occurs during sequential inference. T2M-GPT [47] addresses this issue by introducing random token corruption during training, forcing the model to correct corrupted past contexts. Subsequently, Tian et al. [37] proposed a hierarchical data representation to alleviate cumulative errors. However, these techniques are not directly applicable to conventional autoregressive models—especially diffusion-based architectures with fixed contextual windows [15, 13, 20]. To overcome this limitation, Li et al. [19] proposed a general framework combining an autoregressive context encoder with a lightweight MLP denoiser, leveraging MAE-style random masking [14] to reduce error propagation along the autoregressive direction. Nevertheless, such random masking schemes are less suitable for specially ordered or real-time generation tasks, where temporal consistency is essential. As a result, enhancing the robustness of real-time autoregressive models remains unsolved.
@@ 3 Preliminary
MeanFlow [8] is a recently proposed generative paradigm that aims to produce high-fidelity samples with only a single inference step. Unlike traditional Flow Matching (FM) [22, 25] approaches, which learn the instantaneous velocity field by estimating the continuous trajectory between data and noise distributions through numerous integration steps, MeanFlow instead models the mean velocity over the entire trajectory. Specifically, instead of computing the instantaneous flow $\mathbf{v}(x,t)$ at each time step, MeanFlow evaluates the average transport velocity between any two points $r$ and $t$ in the field. This simplification allows the model to generate samples by a single-step integration of the mean field from $r=0$ (noise) to $t=1$ (data), substantially reducing inference time while maintaining strong generative quality. Our method builds upon MeanFlow, enabling efficient reaction generation with single-step inference. Given an instantaneous velocity field $v(z_{\tau},\tau)$, the average velocity between timesteps $r$ and $t$ is defined as:
@@
where $\mathrm{sg}(\cdot)$ denotes the stop-gradient operator to stabilize optimization. The Jacobian–vector products (JVPs) for $\partial_{z}u_{\theta}$ and $\partial_{t}u_{\theta}$ are computed via automatic differentiation, avoiding explicit Jacobian construction and reducing memory overhead while maintaining accurate gradient signals. Unlike conventional flow-based methods that uniformly sample timestep pairs, we adopt a biased sampling strategy where a proportion of pairs satisfy $r=t$. This allocation strengthens the learning of the instantaneous velocity $v(z_{t},t)$, improves convergence, and enhances fidelity for high-frequency motion components, which are critical for reactive behaviors. A key advantage of MeanFlow is single-step inference, eliminating iterative refinement, which drastically reduces inference cost compared to multi-step diffusion or flow-based models, making it suitable for large-scale generation.
To improve controllability without additional inference overhead, Geng et al. [8] incorporate classifier-free guidance (CFG) during training rather than post-hoc blending by constructing a Ground-Truth Field. The modified objective becomes:
@@ 4 Method
Given a text description texttext and an actor motion sequence $x_{a}\in\mathbb{R}^{T,D}$, where $T$ denotes the number of frames and $D$ the pose dimension, our goal is to synthesize a realistic reactor motion $x_{b}\in\mathbb{R}^{T,D}$. The generation can occur in two modes: sequential (online), where frames are predicted progressively, or offline, where the entire sequence is generated in a single pass. To achieve this, we first compress actor and reactor motions into a shared latent space using the CNN-VAE described in Section 4.1, which benefits both offline and online reaction generation. Building on this representation, we introduce ReMFlow (Section 4.2) for offline generation and ARMFlow (Section 4.3) for online generation, both leveraging the MeanFlow framework and a DiT-based backbone [29]. Finally, we propose the Bootstrap Context Encoding (BCSE), which is inherently suited for ARMFlow and mitigates error accumulation during autoregression.
@@ 4.1 CNN-VAE for Motion Compression
To improve generation efficiency and robustness, we adopt a 1D-CNN VAE following T2M-GPT [47] to compress actor–reactor motion sequences $\mathbf{x}=\{\mathbf{x}_{a},\mathbf{x}_{b}\}$ into temporal latent representations $\{\mathbf{a},\mathbf{b}\}$. This serves to (1) reduce motion dimensionality for efficient computation and (2) provide a structured latent space for autoregressive modeling.
Unlike canonicalized actor motion, reactor motion is highly dynamic and lacks standardization, making discrete quantization prone to accuracy loss. We therefore use continuous latent tokens with KL-divergence regularization instead of discrete codes. This embeds motion into a compact yet expressive latent space, accelerating convergence and improving reconstruction. The causal 1D-CNN preserves temporal dependencies, enabling efficient disentanglement of latent tokens in online generation and enhancing flexibility in sequential contexts. The VAE objective is:
@@
where $q(z|x)$ is the encoder, $p(x|z)$ the decoder, and $p(z)$ the Gaussian prior. To ensure high-fidelity reconstruction, we employ inverse kinematic (IK) loss for joint positions and velocity loss for motion smoothness.
@@ 4.2 ReMFlow for Offline Reaction Generation
We implement ReMFlow for offline generation using a DiT-based encoder with multimodal conditioning. Text prompts are encoded via a CLIP text encoder [26] and fused with timestep embeddings to form a global condition vector, which is injected into DiT through adaptive layer normalization (AdaLN) to modulate intermediate activations. Actor tokens $a_{i}$ are concatenated with interpolated reactor tokens $b_{i}^{t}$, encoded from a CNN-based VAE to capture local appearance and dynamics of the reactor. These tokens are aligned along the last dimension, summed with positional encodings, and processed by DiT to predict average velocities $\mathbf{u}=\{u_{i}\}_{i=1}^{m}$ over tokenized spatiotemporal patches. To enable unconditional learning and support CFG, a proportion of text and actor tokens are replaced with null tokens $\varnothing$ during training, which regularizes the model and stabilizes optimization under varying conditioning strengths.
@@ 4.3 ARMFlow for Online Generation
ARMFlow overcomes the limitations of fixed-window architectures by enabling encoding from scratch and retaining all past information, thereby preserving global semantics. As illustrated in Fig. 2, the architecture follows the spirit of MAR [19], comprising a DiT-based context encoder and a lightweight MLP velocity predictor. The former encodes historical motion and text semantics and supplies them as partial conditions to the MLP predictor.
@@
At inference time, the start token $\langle\text{sos}\rangle$ is first encoded as the initial history. A Gaussian noise token $\epsilon_{1}$ is then paired with the current actor token $a_{1}$ and fed into the MLP with (r=0,t=1)(r{=}0,\,t{=}1) to predict the average velocity $\hat{u}_{1}$. The current reactor token is obtained by $b_{1}=\epsilon_{1}-\hat{u}_{1}$. The predicted reactor token is cached together with the corresponding actor token to update the history, and the process is iterated autoregressively until actor tokens end.
@@ 4.4 Bootstrap Contextual Encoding
Drift over long sequences is an inherent limitation of autoregressive models. HumanX [17] addresses this through a history-rollout training strategy, where the ground-truth (GT) history condition is gradually replaced by the model’s generated history. However, this approach has three main drawbacks. First, to stabilize training, HumanX gradually subtracts the GT history from the generated ones, slowing the convergence. Second, as the model converges, the generated history becomes increasingly similar to the GT, leading to insufficient self-augmentation. Third, HumanX only replaces the reactor history while keeping the actor unchanged, which lead to overfitting to the actor’s motion.
@@
InterFormer [5] $4.871^{\pm.049}$ ${0.302}^{\pm.004}$ ${0.457}^{\pm.004}$ ${0.542}^{\pm.005}$ ${3.845}^{\pm.001}$ $7.482^{\pm.045}$ ${0.254}^{\pm.029}$
CAMDM [3] ${4.000}^{\pm.046}$ ${0.335}^{\pm.005}$ ${0.492}^{\pm.005}$ ${0.587}^{\pm.005}$ ${3.828}^{\pm.001}$ ${7.547}^{\pm.025}$ $\textbf{1.581}^{\pm.026}$
ReGenNet [44] ${4.176}^{\pm.085}$ ${0.355}^{\pm.005}$ ${0.508}^{\pm.005}$ ${0.600}^{\pm.004}$ ${3.817}^{\pm.001}$ ${7.480}^{\pm.033}$ ${0.442}^{\pm.012}$
R2R [1] $\underline{2.795}^{\pm.062}$ $\underline{0.431}^{\pm.005}$ $\underline{0.591}^{\pm.004}$ $\underline{0.674}^{\pm.004}$ $\underline{3.793}^{\pm.002}$ $\underline{7.693}^{\pm.028}$ $\underline{0.517}^{\pm.013}$
@@
InterFormer [5] ${0.304}^{\pm.009}$ ${0.301}^{\pm.003}$ ${0.469}^{\pm.003}$ ${0.571}^{\pm.002}$ ${4.604}^{\pm.009}$ ${8.579}^{\pm.061}$ ${0.289}^{\pm.009}$
CAMDM [3] ${0.429}^{\pm.011}$ ${0.312}^{\pm.004}$ ${0.480}^{\pm.003}$ ${0.587}^{\pm.003}$ ${4.468}^{\pm.020}$ ${8.467}^{\pm.072}$ $\textbf{1.460}^{\pm.027}$
ReGenNet [44] ${0.071}^{\pm.003}$ ${0.402}^{\pm.005}$ ${0.584}^{\pm.004}$ ${0.690}^{\pm.004}$ ${3.843}^{\pm.011}$ $\underline{9.011}^{\pm.053}$ ${0.738}^{\pm.021}$
R2R [1] $\underline{0.063}^{\pm.003}$ $\underline{0.412}^{\pm.005}$ $\underline{0.598}^{\pm.005}$ $\underline{0.704}^{\pm.004}$ $\underline{3.745}^{\pm.011}$ ${8.873}^{\pm.055}$ ${1.074}^{\pm.019}$